# Spectral characterization and performance of SPT-SLIM on-chip filterbank spectrometers

### v2

@@ Abstract
The South Pole Telescope Shirokoff Line Intensity Mapper (SPT-SLIM) experiment is a pathfinder for demonstrating the use of on-chip spectrometers for millimeter (mm) Line Intensity Mapping (LIM). We present spectral bandpass measurements of the SLIM spectrometer channels made on site using a Fourier Transform Spectrometer during SPT-SLIMs first deployment the 2024-2025 austral summer observing season. Through this we demonstrate a technique for measuring the narrow band passes of the SPT-SLIM filterbanks that improves beyond the intrinsic resolution of a Fourier Transform Spectrometer.
@@ I Introduction
Line Intensity Mapping (LIM) uses imaging spectroscopy to trace emission lines as a function of redshift, producing three-dimensional spatial/spectral maps of the large scale structure (LSS) in the Universe without needing to spatially resolve individual sources. In comparison to galaxy surveys, LIM can access larger volumes of space and higher redshifts. LIM in the sub-millimeter/millimeter-wavelength band can provide constraints on cosmology, including expansion history [1] and the neutrino properties, effective number of neutrino species and the sum of the neutrino masses [2]. While it has significant potential, sub-millimeter/millimeter (sub-mm/mm) LIM is still maturing. Significant advancements in the sensitivity and scalability of large field-of-view mapping spectrometers are needed for sub-mm/mm LIM experiments to be competitive with and complementary to modern galaxy surveys.
SPT-SLIM is a pathfinder LIM experiment leveraging integrated on-chip filterbank spectrometers to place constrains on the CO power spectrum in the 120-180 GHz range from the SPT. SPT-SLIM targets CO(2-1) between $0.3<z<0.9$, CO(3-2) between $0.8<z<1.8$, and CO(4-3) between $1.6<z<2.8$ [3]. SPT-SLIM deployed deployed in the austral summer of 2024-2025 on the SPT in the camera position occupied by the Event Horizon Telescope (EHT) very-long-baseline interferometry (VLBI) receiver [4]. SPT is an ideal telescope for mm-wave LIM, as the high-altitude, low humidity, and stable weather at its location limit atmospheric noise [5]. The first deployment of SPT-SLIM has primarily focused on instrumentation and technology demonstration. Bright calibration sources, such as RCW-38 and the Moon, were prioritized to characterize the instrument. Future deployments will have updated detectors and will target a measurement of CO emission lines for LIM science.
@@
As a companion to this work, details on the SPT-SLIM Adiabatic Demagnetization Refrigerator (ADR) used in the deployment are given in Young et al. (2025) [8]. Additionally, details on the RF-ICE detector readout system can be found in Rouble et al. (2022) [9].
@@ II SPT-SLIM: A Filterbank Spectrometer
In the SPT-SLIM design, light is coupled onto polarization sensitive orthomode transducers (OMTs) via gold-plated conical feedhorns [10, 11]. The OMTs are patterned on a thin film and mounted on an aluminum base with $\lambda/4$ backshort features. Each polarization mode of the OMT is coupled via a niobium microstrip transmission line to its own filterbank [12]. Each of the two filterbanks per antenna consist of 65 half-wave resonator filters capacitively coupled to the transmission line. Each filter provides a narrow spectral bandpass at a set frequency of light, with spectral resolution $R=\nu/\Delta\nu$, that then terminates on a Kinetic Inductance Detector (KID). SPT-SLIM uses a Lumped Element KID (LEKID) design with aluminum inductors and niobium interdigitated capacitors. All LEKIDs within a filterbank are coupled to a single microstrip readout line [13]. In addition to the narrow-band detectors that build up the spectrometer, each filterbank has three broadband detectors, and two dark detectors to assist in performance characterization. The dark detectors have no optical coupling to the filterbank transmission line. The SPT-SLIM focal plane consists of three wafers, called “submodules”, each with three readout lines, six filterbanks (one per polarization), and three antennae (see Fig. 1). Due to a combination of damage to the thin film OMTs and non-operational readout lines, five of the nine lines were operational for this first deployment. For additional details on the simulation, design, and fabrication of SPT-SLIM, see Barry et al. (2022) [12], Robson (2024) [13], Robson et al. (2022) [14], and Cecil et al (2023) [15].
SPT-SLIM adds to the burgeoning landscape of on-chip filterbank spectrometers. SPT-SLIM builds on the design and laboratory testing of SuperSpec [16] which has been successfully demonstrated through lab measurements [17] providing a multi-pixel on-sky demonstration of the technique in the 255–278 GHz range. SPT-SLIM also complements the single-pixel DEep Spectroscopic HIgh-redshift MApper (DESHIMA) experiment, operating on-sky in the 220–440 GHz range [18].
@@ III Measurement
Figure 2: The spectral channels of a ploarisation-pair of deployment SPT-SLIM filterbanks measured with the FTS. Note that the width of each filter profile has been broadened by the instrumental line shape of the lower-resolution $\Delta\nu\sim 6$ GHz FTS.
Characterization of the bandpasses of the SPT-SLIM experiment requires both accurate determination of band centers and bandpass shape. An FTS is an interferometer capable of measuring the spectral response of the filterbank spectrometers with an excellent absolute frequency calibration, but with frequency resolution limited by the size of the FTS. A measurement is performed by moving a mirror on a motorized platform at a constant speed, varying the differential path length and introducing a phase difference between a stationary arm and the moving arm of the interferometer. The resulting interference signal (i.e., interferogram) measured by the detectors is the Fourier transform of the detector spectra. Here we note that spectroscopic wavenumbers $\sigma=\frac{1}{\lambda}=\frac{\nu}{c}$, where $c$ is the speed of light and $\nu$ is spectral frequency in units of inverse time, provides a natural frequency unit for FTS measurements. The focal plane was measured with the on-site FTS at the South Pole having a minimum resolution element of $d\nu\sim 6$ GHz, corresponding to $R=25$ at 150 GHz, limited by optical systematics [19]. Additionally, we note that the minimum resolvable spectral resolution element, $\Delta\nu$, is constant across the spectral band.
The SPT-SLIM spectrometers have a design resolution of $R\sim 100$, significantly greater than that of the FTS at the South Pole. As a result, the spectrometer channel bandpasses are unresolved and broadened by the instrumental line shape of the low resolution FTS. Measuring the bandpasses from the spectrum retrieved by the FTS would provide a biased measurement of the filterbank $R$ (see Fig. 2). We demonstrate that an accurate measurement of an under-resolved filter bandpass can be obtained by relying on the assumed spectral profile of the filters. A least-squares fit of the simulated detector signal measured in the spatial displacement domain of the FTS is used to extract the filter resolution and central frequency. For this, we have used the functional form of a Fourier-transformed Lorentzian (the corresponding profile of a half-wave filter),
@@
where $\sigma_{0}$ is the central spectral frequency of the band-pass in wavenumbers and $z$ is optical path difference.
@@
Due to the interaction with other filters in the spectrometer, we expect the Lorentzian band-pass of a single filter to experience some asymmetry [13]. Despite the broadening from the FTS line profile, this can be seen in the measured spectral profile shown in Fig. 3 (lower panel). To account for this, we introduced a skew to the simulated data of the form
@@
where $\tau=R/(\pi\sigma_{0})$ and $\alpha$ is the introduced skew parameter (see Fig. 5). For values ranging $|\alpha|=0$$10\tau$ with a SNR of 10, we have found that the asymmetry does not significantly affect the bias in values of $\sigma_{0}$ and $R$ extracted from spatial-domain fits.
@@
Figure 4: The uncertainty in measurements of the central frequency (top) and spectral resolution/bandpass (bottom) of a filter extracted from spatial-domain fitting using simulated FTS measurements (see Eqs. 13). Simulated measurements were generated with a detector noise model at different signal-to-noise ratios (SNRs) and at different FTS resolving powers. In each case, the ratio of the resolving power of the South Pole FTS at 150 GHz to the $R=100$ simulated filter resolution is marked by a vertical dashed line.
@@ IV Frequency Scheduling and Resolution
Figure 6: The fitted resolution, $R$, of the on-sky submodules extracted from measurements with the South Pole FTS (see Eq. 3). Error bars represent the biases determined from the simulated data shown in Fig. 4 at the corresponding SNR of the interferogram measured with the FTS.
Spatial-domain fitting was applied to the FTS measurements of both SPT-SLIM submodules that were operational for its first deployment at the SPT. All three lines on either submodule were measured, although one was not operational for on-sky measurements. These results are shown in Fig. 6. These measurements show a median spectral resolution of $33\pm 9$ for both deployment submodules, below the design target of $R\approx 100$. We have taken a conservative uncertainty estimate of $\pm 9$ in $R$, corresponding to the bias from the simulated results in Fig. 4 for the mean SNR of the filters in each submodule. Additionally, the centers of the bandpasses are shifted downward from the target SLIM band by $\sim$10 GHz.
Figure 7: A diagram of a single SPT-SLIM channel. Each filter (i.e., half-wave resonator) is capacitive coupled to the transmission line from the OMT and terminates in a capacitive coupled LEKID. Adapted from Robson (2024) [13].
@@
The resolution of a filter is governed by the joined quality factors of the filter’s coupling to the feedline $Q_{\mathrm{feed}}$, the detector coupling, $Q_{\mathrm{det}}$, and losses in the resonator $Q_{\mathrm{loss}}$,
@@
In SPT-SLIM, $Q_{\mathrm{feed}}$ and $Q_{\mathrm{det}}$ are largely controlled by the dimensions of the capacitors that couple the half-wave resonator to the feedline and to the detector, respectively. We expect $Q_{\mathrm{loss}}$ to be dominated by the dielectric loss tangent ($\tan\delta$) of the SiN employed in the microstrip architecture.
@@
The dielectric loss tangent of the SPT-SLIM structures compromises its performance in nearly every aspect, reducing optical efficiency, resolution, and increasing optical loading from the atmospheric emission lines that bookend the SLIM band.
@@ V Conclusion
In the 2024-2025 austral summer, SPT-SLIM deployed on the SPT in the camera position occupied by the EHT VLBI receiver. This deployment included 5 operational dual-polarization filterbank spectrometers. These spectrometers were characterized on site with a low-resolution FTS. We demonstrate a technique by which the spectral bandpass/resolution of the filters can be measured with an FTS having less than a quarter the resolution of the designed filter bandwidth. With this, the SPT-SLIM filterbanks deployed were measured to have an average spectral resolution $R=34$, lower than the target resolution of $R=100$. This is likely caused by differences between the simulated and fabricated dielectrics, in which lossy dielectric lowers the resolution from design. These measured bandpasses are applied to observations of the atmosphere and moon in Dibert et al. [22].
@@ VI Acknowledgments
This work was supported by funding from the National Science Foundation, Fermilab National Laboratory under the Department of Energy, and the United Kingdom Research and Innovation.

### v3

@@ Abstract
The South Pole Telescope Shirokoff Line Intensity Mapper (SPT-SLIM) experiment is a pathfinder for demonstrating the use of on-chip spectrometers for millimeter (mm) Line Intensity Mapping (LIM). We present spectral bandpass measurements of the SLIM spectrometer channels made on site using a Fourier Transform Spectrometer during the first deployment of SPT-SLIM in the 2024-2025 Austral summer season. Through this, we demonstrate a technique for measuring the narrow band passes of the SPT-SLIM filterbanks that improves beyond the intrinsic resolution of a Fourier Transform Spectrometer.
@@ I Introduction
Line Intensity Mapping (LIM) uses imaging spectroscopy to trace emission lines as a function of redshift, producing three-dimensional spatial/spectral maps of the large scale structure (LSS) in the Universe without needing to spatially resolve individual sources. In comparison to galaxy surveys, LIM can access larger volumes of space and higher redshifts. LIM in the sub-millimeter/millimeter-wavelength band can provide constraints on cosmology, including expansion history [1] and the neutrino properties, effective number of neutrino species and the sum of the neutrino masses [2]. While it has significant scientific potential, sub-millimeter/millimeter (sub-mm/mm) LIM is still maturing. Significant advancements in the sensitivity and scalability of large field-of-view mapping spectrometers are needed for sub-mm/mm LIM experiments to be competitive with and complementary to modern galaxy surveys.
SPT-SLIM is a pathfinder LIM experiment leveraging integrated on-chip filterbank spectrometers to place constraints on the CO power spectrum in the 120-180 GHz range from the SPT. SPT-SLIM targets CO(2-1) between $0.3<z<0.9$, CO(3-2) between $0.8<z<1.8$, and CO(4-3) between $1.6<z<2.8$ [3]. SPT-SLIM deployed in the Austral summer of 2024-2025 on the SPT in the camera position previously occupied by the Event Horizon Telescope (EHT) very-long-baseline interferometry (VLBI) receiver [4]. The SPT is an ideal telescope for mm-wave LIM, as the high altitude, low humidity, and stable weather at its location limit atmospheric noise [5]. The first deployment of SPT-SLIM has primarily focused on instrumentation and technology demonstration. Bright calibration sources, such as RCW-38 and the Moon, were prioritized to characterize the instrument. Future deployments will have updated detectors and will target a measurement of CO emission lines for LIM science.
@@
As a companion to this work, details on the SPT-SLIM Adiabatic Demagnetization Refrigerator (ADR) used in the deployment are given in Young et al. (2025) [8] and the atmospheric calibration of the instrument is outlined in Karia et al (2025) [9]. Additionally, details on the RF-ICE detector readout system can be found in Rouble et al. (2022) [10].
@@ II SPT-SLIM: A Filterbank Spectrometer
In the SPT-SLIM design, light is coupled onto polarization-sensitive orthomode transducers (OMTs) via gold-plated conical feedhorns [11, 12]. The OMTs are patterned on a thin film and mounted on an aluminum base with $\lambda/4$ backshort features. Each polarization mode of the OMT is coupled via a niobium microstrip transmission line to its own filterbank [13].
Each of the two filterbanks per antenna consists of 65 half-wave resonator filters capacitively coupled to the transmission line. Each filter provides a narrow spectral bandpass at a set frequency of light, with spectral resolution $R=\nu/\Delta\nu$, that then terminates on a Kinetic Inductance Detector (KID). SPT-SLIM uses a Lumped Element KID (LEKID) design with aluminum inductors and niobium interdigitated capacitors. All LEKIDs within a filterbank are coupled to a single microstrip readout line [14]. In addition to the narrow-band detectors that build up the spectrometer, each filterbank has three broadband detectors, and two dark detectors to assist in performance characterization. The dark detectors have no optical coupling to the filterbank transmission line.
The SPT-SLIM focal plane consists of three wafers, called “submodules”, each with three readout lines, six filterbanks (one per polarization), and three antennae (see Fig. 1). Due to a combination of damage to the thin film OMTs and non-operational readout lines, five of the nine lines were operational for this first deployment. For additional details on the simulation, design, and fabrication of SPT-SLIM, see Barry et al. (2022) [13], Robson (2024) [14], Robson et al. (2022) [15], and Cecil et al (2023). [16].
SPT-SLIM adds to the burgeoning landscape of on-chip filterbank spectrometers providing a multi-pixel on-sky demonstration of the technique. Additionally, SPT-SLIM builds on the design and laboratory testing of SuperSpec [17] which has been successfully demonstrated through lab measurements in the 255–278 GHz range [18]. SPT-SLIM also complements the single-pixel DEep Spectroscopic HIgh-redshift MApper (DESHIMA) experiment, operating on-sky in the 220–440 GHz range [19].
@@ III Measurement
The spectral channels of a polarization-pair of deployment SPT-SLIM filterbanks measured with the FTS. Note that the width of each filter profile has been broadened by the instrumental line shape of the lower-resolution $\Delta\nu\sim 6$ GHz FTS.
Characterization of the bandpasses of the SPT-SLIM experiment requires both accurate determination of band centers and bandpass shape. A Fourier Transform Spectrometer (FTS) is an interferometer capable of measuring the spectral response of the filterbank spectrometers with an excellent absolute frequency calibration, with frequency resolution ultimately limited by the size of the FTS. A measurement is performed by moving a mirror on a motorized platform at a constant speed, varying the differential path length and introducing a phase difference between a stationary arm and the moving arm of the interferometer. The resulting interference signal (i.e., interferogram) measured by the detectors is the Fourier transform of the detector spectra. Here we note that spectroscopic wavenumbers $\sigma=\frac{1}{\lambda}=\frac{\nu}{c}$, where $c$ is the speed of light and $\nu$ is spectral frequency in units of inverse time, provide a natural frequency unit for FTS measurements. The focal plane was measured with the on-site FTS at the South Pole having a minimum resolution element of $d\nu\sim 6$ GHz, corresponding to $R=25$ at 150 GHz, primarily limited by optical systematics [20, 21]. Additionally, we note that the minimum resolvable spectral resolution element of an FTS, $\Delta\nu$, is constant across the spectral band. Following from the simulation work we outline in this work, we deemed the low resolution of this FTS that was available on site sufficient for an initial characterization of the SLIM spectrometers while following up with a higher-resolution measurement post-deployment in a laboratory setting.
The SPT-SLIM spectrometers have a design resolution of $R\sim 100$, significantly greater than that of the FTS at the South Pole. As a result, the spectrometer channel bandpasses are unresolved and broadened by the instrumental line shape of the low-resolution FTS. Measuring the bandpasses from the spectrum retrieved by the FTS would provide a biased measurement of the filterbank $R$ (see Fig. III).
We demonstrate that an accurate measurement of an under-resolved filter bandpass can be obtained by relying on the assumed spectral profile of the filters. Using a least-squares fit of the detector signal measured in the spatial displacement domain of the FTS, we extract the filter resolution and central frequency. For this, we have used the functional form of a Fourier-transformed Lorentzian (the corresponding profile of a half-wave filter),
@@
where $\sigma_{0}$ is the central spectral frequency of the bandpass in wavenumbers and $z$ is optical path difference.
@@
Due to the interaction with other filters in the spectrometer, we expect the Lorentzian bandpass of a single filter to experience some asymmetry [14]. Despite the broadening from the FTS line profile, this can be seen in the measured spectral profile shown in Fig. 2 (lower panel). To account for this, we introduced a skew to the simulated data of the form
@@
where $\tau=R/(\pi\sigma_{0})$ and $\alpha$ is the introduced skew parameter (see Fig. 4). For values ranging $|\alpha|=0$$10\tau$ with a SNR of 10, we have found that the asymmetry does not significantly affect the values of $\sigma_{0}$ and $R$ extracted from spatial-domain fits.
@@
Figure 3: The disagreement between simulated values and measurements of the central frequency (top) and spectral resolution/bandpass (bottom) of a filter extracted from spatial-domain fitting using simulated FTS measurements (see Eqs. 13). Simulated measurements were generated with a detector noise model at different signal-to-noise ratios (SNRs) and at different FTS resolving powers with each bin/pixel showing the mean of 20 simulated measurements. In each panel, the ratio of the resolving power of the South Pole FTS at 150 GHz to the $R=100$ simulated filter resolution is marked by a vertical dashed line. The contours in the bottom figure mark uncertainty thresholds in values of $R$ determined from this method determined from within the parameter space.
@@ IV Frequency Scheduling and Resolution
Figure 5: The fitted resolution, $R$, of the on-sky submodules extracted from measurements with the South Pole FTS (see Eq. 3). Error bars represent the uncertainty determined from the simulated data shown in Fig. 3 at the corresponding SNR of the interferogram measured with the FTS.
Spatial-domain fitting was applied to the FTS measurements of both SPT-SLIM submodules that were operational for its first deployment at the SPT. All three lines on either submodule were measured, although one was not operational for on-sky measurements. These results are shown in Fig. 5. These show a median spectral resolution of $33\pm 9$ for both deployment submodules, below the design target of $R\approx 100$. We have taken a conservative uncertainty estimate of $\pm 9$ in $R$, corresponding to the error from the simulated results in Fig. 3 for the mean SNR of the filters in each submodule. Additionally, the centers of the bandpasses are shifted downward from the target SLIM band by $\sim$10 GHz.
Figure 6: A diagram of a single SPT-SLIM channel. Each filter (i.e., half-wave resonator) is capacitively coupled to the transmission line from the OMT and terminates in a capacitive coupled LEKID. Adapted from Robson (2024) [14].
@@
The resolution of a filter is governed by the joined quality factors of the filter’s coupling to the feedline, $Q_{\mathrm{feed}}$, the detector coupling, $Q_{\mathrm{det}}$, and losses in the resonator, $Q_{\mathrm{loss}}$,
@@
In SPT-SLIM, $Q_{\mathrm{feed}}$ and $Q_{\mathrm{det}}$ are largely controlled by the dimensions of the coupling capacitors of the half-wave resonator to the feedline and to the detector, respectively. We expect $Q_{\mathrm{loss}}$ to be dominated by the dielectric loss tangent ($\tan\delta$) of the SiN employed in the microstrip architecture.
@@
The dielectric loss tangent of the SPT-SLIM structures compromises its performance in nearly every aspect, reducing optical efficiency, resolution, and increasing optical loading from the atmospheric emission lines that bookend the SLIM band. Lower spectral resolution reduces the resolution in redshift, and therefore reduces potential constraints on cosmological parameters. We expect significant improvements in this respect as we explore alternative low-loss dielectrics while we continue to iterate on our fabrication processes and expand our fabrication potential.
@@ V Conclusion
In the 2024-2025 Austral summer, SPT-SLIM deployed on the SPT in the camera position occupied by the EHT VLBI receiver. This deployment included 5 operational dual-polarization filterbank spectrometers. These spectrometers were characterized on site with a low-resolution FTS. We demonstrate a technique by which the spectral bandpass/resolution of the filters can be measured with an FTS having less than a quarter the resolution of the designed filter bandwidth. With this, the SPT-SLIM filterbanks deployed were measured to have an average spectral resolution $R=34$, lower than the target resolution of $R=100$. This is likely caused by differences between the simulated and fabricated performance of the SiN dielectric and coupling capacitors that have been employed. These measured bandpasses are applied to observations of the atmosphere and moon in Dibert et al. [9].
@@ VI Acknowledgments
This work was supported by funding from the National Science Foundation (NSF), Fermilab National Laboratory under the Department of Energy, and the United Kingdom Research and Innovation (UKRI). SPT-SLIM is supported by the National Science Foundation under Award No. AST-2108763. The South Pole Telescope program is supported by the NSF through awards OPP-1852617 and OPP-2332483. Work at Argonne, including use of the Center for Nanoscale Materials, an Office of Science user facility, was supported by the US Department of Energy, Office of Science, Office of Basic Energy Sciences and Office of High Energy Physics, under Contract No. DE-AC02-06CH11357. This document was prepared by the SPT-SLIM collaboration using the resources of the Fermi National Accelerator Laboratory (Fermilab), a U.S. Department of Energy, Office of Science, Office of High Energy Physics HEP User Facility. Fermilab is managed by FermiForward Discovery Group, LLC, acting under Contract No. 89243024CSC000002. The McGill team acknowledges funding from the Natural Sciences and Engineering Research Council of Canada and the Canadian Institute for Advanced Research, and the Canada Research Chairs program. This work is supported by UKRI Future Leaders Fellowship MR/W006499/1. Partial support is also provided by the Kavli Institute of Cosmological Physics at the University of Chicago. Support for this work for JZ was provided by NASA through the NASA Hubble Fellowship grant HF2-51500 awarded by the Space Telescope Science Institute, which is operated by the Association of Universities for Research in Astronomy, Inc., for NASA, under contract NAS5-26555. KSK was partially supported by an NSF Astronomy and Astrophysics Postdoctoral Fellowship under award AST-2001802, and by SLAC under award LDRD-24-008. We thank Heitor Mourato, Glenn Thayer, and Jose Velho in the BU Scientific Instrument Facility for assistance with mirror fabrication, and John Kovac and Miranda Eiben for assistance with anti-reflection coating optics.